Agenda/Minutes on 2020/12/4
CRY+VIS Interface Meeting
Participants: Takahashi, Hagiwara, Yano, Miyo, Ikeda, Sato, Ushiba, Aso, Kimura, Hirata
- Schedule: Upgrade work in IXV (2/22-3/10) will be swapped with the work in EXV (1/4-1/22). The connection of ITMX is available on 1/12.
- Jig for the FC painting.
Acceptance check items (Inspection document).
PF actuator: Measurement of coefficients --> calculation of the noise (by the end of Jan.).
Participants: Takahashi, Yano, Miyo, Ikeda, Sato, Aso, Fabian, Hirata, Iguchi, Tanaka, Terrence
As you all know, we have had trouble starting a measurement with a DGS error klog#15567. In order to eliminate this confusion during the work, I make the following rules;
First, make sure that the VIS_OVERVIEW screen shows all green before you start working.
And after working, make everything green before finishing your work. This is a kindness for the next worker.
The SDF will probably turn red due to your work. This means that SDF has detected EPICS channel differences during your work. If you accept these differences, please take a screenshot and post it in klog (e.g. klog#15553), but you don't have to post when you revert.
- That's all.
[Fabian] The thickness in the cables we are using for the stepper motors seems fine. Could Yano-san and Hirata-san explain, please?
We inspected the IP LVDT-actuator units and horizontal fishing rods. See pictures in this album.
- The yokes have some rust and coils have dust.
- The horizontal FR blades don't have nickel plating but they are not rusty in an obvious way.
- We found the jigs to remove the LVDTs.
- We checked there's suitable access to install Satou-san's limit switches.
- Terrence did various measurements and we locked and unlock the IP for it.
- Temperature in the clean booth is high (24.7 degrees). I kindly asked Yokozawa-san to take action to cool it down.
SRM, SR3 and SR2
- We finished the installation of the F0 FR limit switches after we got new fasteners.
- In SRM and SR2 we compensated the additional mass of the yaw stepper motor assemblyby removing ballast mass from the keystone.
- In SR3 we set the F0 FR upper limit switch to a suitable height.
- We tested the switches by pressing them manually and checking the signal in the medm screen using the update button. They worked fine this way.
- However, some of the Nikhef switches in the IPs require fixing. They are made of copper sheet and they are bent in such a way that system receives a trigger signal even when they are not pressed. The cales are fine because they react when they are pushed by hand.
- SR2: H1, -Yaw, Left; H2, -Yaw, Left.
- SR3: H1, -Yaw, Left.
- SRM: We fixed one.
- Tthe script doesn't automatically warns the human user when a switch has been pressed. Currently, the update button has to be pressed to know the status. Ikeda-san will implement an automatic warning by the end of January.
Ikeda-san clarified the driver should stop the motor as soon as the switch is pressed. We did an experiment that is shown in this short movie (SR3).
- In SR3 we connected again the satellite boxes to the rack.
A TO DO list for SR suspensions.
- Fix the IP limits switches which are not working.
Recover the local alignment of the suspensions and measure transfer functions.
- Change polarity of SRM F0 FR stepper motor.
- In SR3 exchange the heavy top plate (1 kg) from the keystone assembly by a lighter new one (less than 400 g).
- Set all the switches to suitable positions.
- Make sure the script automatically detects the pressing of the switches.
- Repair connector adapter for picomotors. Check whether there are missing ones.
- Connect BF picomotors in BS and SR3.
- SRM IM-R picomotor is not working. Check it.
- Decide what to do about the position of the F0 keystones: they are very low because we lifted the suspensions with the jacks in order to compensate for the buoyancy effect. However, Mark made us change our minds and we didn't lowered them back before O3GK.
- [Takahashi, Iguchi]
- Set up the old OSTM suspension in the OMC both.
I started working with Ikeda-san and Yamamoto-san to standardize the VIS models. Purpose of this work is to eliminates subtle model differences, such as different naming conventions for different suspension types.
By doing so, we will be able to do the same commissioning if the hardware is the same, so that more suspensions can be commissioned by fewer people.
- In this week, we are working on arranging the models of the Tower except the GAS. I'll get consensus from Terrence and Fabian about changing the models.
- [Takahashi] Tested the modified LVDT card. This modification (#14 pin grounded) will be applied to 16 cards.
- [Tanaka] The prototype circuit was completed and under the measurement. The performance at the frequencies lower than 100Hz is not good. The attenuation is too small. Still measuring.
- Type-B remedy work is on going.
- Type-A team will start modification of ACCs.
Upgrade work in IXV (2/22-3/10) will be swapped with the work in EXV (1/4-1/22). (Chart)
- Procedure documents.
- Criteria in the acceptance test.
- Mechanical resonances of SRM.
- Life time of the LDs for Oplev. It is necessary to turn on-off frequently.
- When the Oplevs should be return?
- Information of the line frequencies in Type-B.
The week from 12/7
- Takahashi: 12/7-11
- Hirata: 12/7-11
- Yano: 12/7-11
- Sato: 12/7-11
The week from 12/14
- Takahashi: 12/14-18
- Hirata: 12/14-18
- Yano: 12/14-18
(F2F Meeting in 12/17-18)
- On 2020/12/11(Fri)